# Exceptionally High Carrier Mobility in Hexagonal Diamond

### v1

@@ Abstract
For electron scattering, the selection rules along three Cartesian directions are evaluated. In c-diamond, the constant-energy surfaces near the CBM [at , with a Cartesian coordinate of ] manifest as ellipsoids. The selection rule for intravalley scattering along the principal axis is
@@
where the Cartesian coordinate of the final state is . In h-diamond (CBM at ), the selection rules are
@@
In both phases, LA phonons (, , and symmetry in h-diamond) are allowed to induce scattering, whereas transitions mediated by TA modes (, , , , and symmetry in h-diamond) are forbidden in specific directions. Compared with c-diamond, TA phonons in h-diamond are subject to stricter restrictions, which reduces the number of scattering channels, similar to the case of hole transport. Consequently, the TA-limited electron mobility in c-diamond is approximately 2/3 of its actual mobility at 300 K, whereas in h-diamond, this ratio is below 30%. Nevertheless, the effect of optical phonons is more significant in h-diamond than in c-diamond, which partially offsets that of the suppressed TA-scattering. Therefore, the difference in the selection rules may be insufficient to account for the substantially higher electron mobility in h-diamond. Instead, the charge distribution provides a possible explanation. As shown in Fig. 3(b), the CBM state in h-diamond is composed of antibonding and orbitals, consistent with the point symmetry at the points. The corresponding charge density exhibits an extended distribution within the interstitial regions of the lattice, giving rise to a small effective mass. Moreover, scattering potentials are primarily concentrated near the atoms and bonds (as nonpolar h-diamond lacks long-range Fröhlich interaction), leading to weak overlap with the electronic wavefunction at CBM, as shown in Fig. S8 in SM [1]. Such a “de-coupling” between electrons and phonons considerably decreases the values of matrix elements (Fig. S9 in SM [1]), thereby reducing the scattering rate, whereby playing an important role in the exceptionally high electron mobility of h-diamond.
@@
In summary, the carrier mobilities in h-diamond and c-diamond are calculated using ab initio methods. At 300 K, the hole mobility perpendicular and parallel to the -axis is 5631 and 5552 cm2V-1s-1 in h-diamond, respectively, and the electron mobility along the two directions is 11462 and 28464 cm2V-1s-1. These values are much higher than the mobility of c-diamond, in which the room-temperature hole and electron mobilities are 2543 and 2068 cm2V-1s-1, respectively. In c-diamond, the ADP scattering mediated by TA phonons is the predominant scattering mechanism at room temperature, which is considerably suppressed in h-diamond by the selection rules. Although the interlayer shear and breathing modes provide additional scattering channels in h-diamond, they have weak interactions with the VBM state, which exhibits an in-plane distribution. At the CBM, the wavefunction of the CBM state is composed of antibonding and orbitals, which is extended and distributed mainly in the interstitial regions within the lattice, giving rise to spatial “de-coupling” with scattering potentials, resulting in high electron mobility. Other factors including effective mass have little contribution to the difference in carrier mobility between the two phases. We also show the temperature dependence of carrier mobility. At elevated temperatures, with the enhanced occupations of high-energy phonons, optical deformation potential scattering becomes the predominant mechanism, leading to a deviation from the power-law relationship. Our findings indicate that h-diamond is a promising candidate for high-frequency and high-speed applications. Furthermore, this work also offers a comprehensive picture of the carrier transport property of semiconductors. Although simple parameters, particularly effective mass, can usually provide a basic and straightforward understanding of carrier mobility, other complicated factors, such as the symmetry and orbital component of the band extremum point, may also play a considerable role.
@@ Acknowledgements.
- [1] Note: http://link.aps.org/supplemental/10.1103/PhysRevB.XX.XXXXXX Cited by: Table 1, Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [2] (1982-10) Semiconducting and other major properties of gallium arsenide. J. Appl. Phys. 53 (10), pp. R123–R181. External Links: ISSN 0021-8979, Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [3] (1967-05) Hexagonal diamond—a new form of carbon. J. Chem. Phys. 46 (9), pp. 3437–3446. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [4] (2018) Lattice thermal transport in superhard hexagonal diamond and wurtzite boron nitride: a comparative study with cubic diamond and cubic boron nitride. Carbon 139, pp. 85–93. External Links: Link, Document Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [5] (2008) Group theory: application to the physics of condensed matter. Springer Berlin, Heidelberg. External Links: Document Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [6] (1967) Lonsdaleite, a hexagonal polymorph of diamond. Nature (London) 214 (5088), pp. 587–589. External Links: Document Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [7] (2011-03) Electron and hole drift velocity in chemical vapor deposition diamond. J. Appl. Phys. 109 (6), pp. 063719. External Links: ISSN 0021-8979, Document, Link Cited by: Figure 1.
@@
- [8] (2018-10) -Ga2O3 for wide-bandgap electronics and optoelectronics. Semicond. Sci. Technol. 33 (11), pp. 113001. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [9] (1967) Hexagonal diamonds in meteorites: implications. Science 155 (3765), pp. 995–997. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [10] (1969) High purity GaAs by liquid phase epitaxy. Solid State Commun. 7 (20), pp. 1463–1465. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [11] (2018) Experimental observation of high thermal conductivity in boron arsenide. Science 361 (6402), pp. 575–578. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [12] (2018) High thermal conductivity in cubic boron arsenide crystals. Science 361 (6402), pp. 579–581. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [14] (1954-01) Lattice-scattering mobility in germanium. Phys. Rev. 93, pp. 62–63. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [15] (2009-02) Harder than diamond: superior indentation strength of wurtzite BN and lonsdaleite. Phys. Rev. Lett. 102, pp. 055503. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [16] (2021-10) First-principles predictions of Hall and drift mobilities in semiconductors. Phys. Rev. Res. 3, pp. 043022. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [17] (2021) Phonon thermal transport in diamond and lonsdaleite: a comparative study of empirical potentials. Diam. Relat. Mater. 120, pp. 108618. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [18] (2016-05) Single crystal diamond wafers for high power electronics. Diam. Relat. Mater. 65 (SI), pp. 168–175. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [19] (2022-JUL 22) High ambipolar mobility in cubic boron arsenide. Science 377 (6604), pp. 437–440. External Links: Document, ISSN 0036-8075 Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [20] (2018) Unusual high thermal conductivity in boron arsenide bulk crystals. Science 361 (6402), pp. 582–585. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [21] (2008) Diamond as an electronic material. Mater. Today 11 (1-2), pp. 22–28. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [22] (2025) Synthesis of bulk hexagonal diamond. Nature (London) 644 (8076), pp. 370–375. External Links: Link, Document Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [23] (2024-10) First-principles investigation of microscopic mechanisms underlying hole mobilities in diamond, silicon, and germanium. Phys. Rev. B 110, pp. 155203. External Links: Document, Link Cited by: footnote 1, Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [24] (2022-JUL 22) High ambipolar mobility in cubic boron arsenide revealed by transient reflectivity microscopy. Science 377 (6604), pp. 433–436. External Links: Document, ISSN 0036-8075 Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [25] (2025-09) Thermal properties of hexagonal diamond: machine learning potential and molecular dynamics study. Phys. Rev. Mater. 9, pp. 094603. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond, Exceptionally High Carrier Mobility in Hexagonal Diamond.
@@
- [26] (2024) Pure hexagonal diamond with symmetry-doping properties. ACS Appl. Mater. Interfaces 16 (39), pp. 52674–52681. External Links: Document, Link Cited by: Exceptionally High Carrier Mobility in Hexagonal Diamond.

### v2

@@ II.4 Origin of the higher electron mobility in h-diamond
For electron scattering, the selection rules along the three Cartesian directions are evaluated. In c-diamond, the constant-energy surfaces near the CBM [located at the point with a Cartesian coordinate of ] are ellipsoids. The selection rule for intravalley scattering along the principal axis is
@@
where the point has a Cartesian coordinate of . In h-diamond (CBM at ), the selection rules are
In both phases, LA phonons (of , , and symmetry in h-diamond) are allowed to induce scattering, whereas transitions mediated by TA modes (of , , , , and symmetry in h-diamond) are forbidden in specific directions. Compared with c-diamond, TA phonons in h-diamond are subject to stricter restrictions, which reduce the number of scattering channels, similar to the case of hole transport. Consequently, the TA modes contribute about 2/3 of the total scattering in c-diamond at 300 K, whereas such a contribution is below 30% in h-diamond. Nevertheless, optical phonons have a more significant influence in h-diamond than in c-diamond, partially offsetting the effect of suppressed TA-scattering. Therefore, the difference in selection rules may be insufficient to account for the substantially higher electron mobility in h-diamond.
@@
Instead, the charge distribution provides a possible explanation. As shown in Fig. 4(b), the CBM state in h-diamond is composed of antibonding - orbitals, consistent with the point-group symmetry at the points. The corresponding charge density is delocalised over the interstitial regions of the lattice, contributing to a small effective mass. Moreover, since h-diamond is nonpolar and thus lacks the long-range Fröhlich interaction, scattering potentials are concentrated near atoms and bonds, leading to weak spatial overlap with the CBM wavefunction, as visualised in Fig. 5. Such a “decoupling” between electrons and phonons considerably decreases the values of matrix elements (Fig. S8 of the SI), thereby reducing the scattering rate and playing a crucial role in the exceptionally high electron mobility of h-diamond.
III Discussion
@@
In summary, we reveal the exceptionally high mobility in h-diamond and elucidate its physical origins. ADP scattering mediated by TA phonons, which is the predominant scattering mechanism for holes in c-diamond around room temperature, is considerably suppressed in h-diamond by selection rules. The in-plane distribution of the VBM wavefunction also leads to weak interactions with the additional interlayer shear and breathing modes in h-diamond. For electrons, the CBM wavefunction formed by antibonding - orbitals is distributed primarily in the interstitial regions within the h-diamond lattice, giving rise to little overlap with scattering potentials, thus significantly enhancing electron mobility. Other factors including effective mass are comparable in the two phases, contributing negligibly to their mobility difference. We also show the temperature dependence of carrier mobility, which deviates from the power-law relationship due to the increased occupation of high-energy optical phonons at elevated temperatures. Our findings indicate h-diamond as a promising candidate for high-frequency and high-speed applications. Furthermore, by revealing the significance of selection rules and electron-phonon “decoupling”, this work offers fresh insights into the mechanisms associated with high carrier mobility.
IV Methods
IV.1 DFT and DFPT calculations
Geometry optimisation and phonon calculations were performed using the quantum espresso software [10, 8, 9]. The Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm was utilised for geometry optimisation, with the convergence thresholds for energy, force, and stress set to Ry, Ry/Bohr, and 0.05 kbar, respectively. Phonon calculations were based on density-functional perturbation theory (DFPT), with a convergence threshold of . Optimised norm-conserving Vanderbilt pseudopotentials [12] with full relativistic treatment from the pseudo dojo project [35] were used, and the plane-wave cutoff energy was chosen to be 100 Ry. The PBEsol functional [24] was applied for the approximation of exchange-correlation interaction.
IV.2 band structure calculations
In order to obtain accurate - dispersion relationships, quasi-particle eigenvalues were calculated using the one-shot method, as implemented in the yambo software [19, 29]. The band structures and DOS were then obtained based on the maximally localised Wannier functions (MLWFs) [20, 33] as implemented in the wannier90 code [25].
IV.3 e-ph scattering and mobility calculations
For the calculations of e-ph scattering and carrier mobility, the electronic Hamiltonians, dynamical matrices, and electron-phonon matrix elements were interpolated onto ultrafine and grids based on MLWFs using the epw code [26, 17]. The energy window was set to 0.3 eV (relative to the band edge), and the grid sizes are listed in Table S2 of the SI. The spin-orbit coupling (SOC) was found to have a negligible effect on electron mobility and was thus only considered for valence bands. To ensure the accurate description of long-range interactions, the effect of dynamical quadrupoles was included, which were calculated based on DFPT using the abinit software [11, 28].
IV.4 MD simulation
In order to examine the thermal stability of h-diamond, we performed a MD simulation in the ensemble at 1000 K and 1 atm. The time step was chosen to be 1 fs, and the MD simulation ran for 10,000 steps (i.e., 10 ps). Forces and pressures were calculated based on DFT using the Gaussian and plane wave (GPW) method as implemented in the the cp2k code [16], employing the DZVP-MOLOPT basis set. The cutoff energy and the relative cutoff energy were set to 300 and 50 Ry, respectively. The exchange-correlation interaction was approximated by the PBEsol functional [24]. A supercell containing 1280 atoms was used, and the BZ was sampled with a single -point.
Data availability
The data that support the findings of this work are available from the corresponding authors upon reasonable requests.
Code availability
All computational codes used in this work (quantum espresso, epw, abinit, yambo, wannier90, and cp2k) are open-source and available online.
Competing interests
The authors declare no competing interests.
Author contributions
Z. H., S. G., and M. C. conceptualised and designed this study. Z. H. performed the calculations, analysed the results, and wrote the initial draft. S. G. and M. C. edited the manuscript and supervised the project. All authors reviewed and approved the final version of the manuscript.
@@ Acknowledgements.
- [1] (1982-10) Semiconducting and other major properties of gallium arsenide. J. Appl. Phys. 53 (10), pp. R123–R181. External Links: ISSN 0021-8979, Document, Link Cited by: §I.
@@
- [2] (1967-05) Hexagonal diamond—a new form of carbon. J. Chem. Phys. 46 (9), pp. 3437–3446. External Links: Document, Link Cited by: §I.
@@
- [3] (2018) Lattice thermal transport in superhard hexagonal diamond and wurtzite boron nitride: a comparative study with cubic diamond and cubic boron nitride. Carbon 139, pp. 85–93. External Links: Link, Document Cited by: §I, §II.1.
@@
- [4] (2008) Group theory: application to the physics of condensed matter. Springer Berlin, Heidelberg. External Links: Document Cited by: §II.3.
@@
- [5] (1967) Lonsdaleite, a hexagonal polymorph of diamond. Nature (London) 214 (5088), pp. 587–589. External Links: Document Cited by: §I.
@@
- [6] (2011-03) Electron and hole drift velocity in chemical vapor deposition diamond. J. Appl. Phys. 109 (6), pp. 063719. External Links: ISSN 0021-8979, Document, Link Cited by: Figure 1, §II.1.
@@
- [7] (2018-10) -Ga2O3 for wide-bandgap electronics and optoelectronics. Semicond. Sci. Technol. 33 (11), pp. 113001. External Links: Document, Link Cited by: §I.
- [8] (2017-10) Advanced capabilities for materials modelling with Quantum ESPRESSO. J. Phys.: Condens. Matter 29 (46), pp. 465901. External Links: Document, Link Cited by: §IV.1.
- [9] (2020-04) Quantum ESPRESSO toward the exascale. J. Chem. Phys. 152 (15), pp. 154105. External Links: ISSN 0021-9606, Document, Link Cited by: §IV.1.
- [10] (2009-09) QUANTUM ESPRESSO: a modular and open-source software project for quantum simulations of materials. J. Phys.: Condens. Matter 21 (39), pp. 395502. External Links: Document, Link Cited by: §IV.1.
- [11] (2020) The abinit project: impact, environment and recent developments. Comput. Phys. Commun. 248, pp. 107042. External Links: Link Cited by: §IV.3, §IV.3.
- [12] (2013-08) Optimized norm-conserving Vanderbilt pseudopotentials. Phys. Rev. B 88, pp. 085117. External Links: Document, Link Cited by: §IV.1.
@@
- [13] (1967) Hexagonal diamonds in meteorites: implications. Science 155 (3765), pp. 995–997. External Links: Document, Link Cited by: §I.
@@
- [14] (1969) High purity GaAs by liquid phase epitaxy. Solid State Commun. 7 (20), pp. 1463–1465. External Links: Document, Link Cited by: §II.1.
@@
- [15] (2018) Experimental observation of high thermal conductivity in boron arsenide. Science 361 (6402), pp. 575–578. External Links: Document, Link Cited by: §II.1.
- [16] (2020-05) CP2K: an electronic structure and molecular dynamics software package - quickstep: efficient and accurate electronic structure calculations. J. Chem. Phys. 152 (19), pp. 194103. External Links: ISSN 0021-9606, Document, Link Cited by: §IV.4.
- [17] (2023) Electron–phonon physics from first principles using the EPW code. npj Comput. Mater. 9, pp. 156. External Links: Document Cited by: §IV.3.
@@
- [18] (2018) High thermal conductivity in cubic boron arsenide crystals. Science 361 (6402), pp. 579–581. External Links: Document, Link Cited by: §II.1.
- [19] (2009-08) Yambo: an ab initio tool for excited state calculations. Comp. Phys. Commun. 180 (8), pp. 1392–1403. External Links: Document, ISSN 0010-4655 Cited by: §IV.2.
- [20] (1997-11) Maximally localized generalized Wannier functions for composite energy bands. Phys. Rev. B 56, pp. 12847–12865. External Links: Document, Link Cited by: §IV.2.
@@
- [22] (1954-01) Lattice-scattering mobility in germanium. Phys. Rev. 93, pp. 62–63. External Links: Document, Link Cited by: §II.1.
@@
- [23] (2009-02) Harder than diamond: superior indentation strength of wurtzite BN and lonsdaleite. Phys. Rev. Lett. 102, pp. 055503. External Links: Document, Link Cited by: §I.
- [24] (2008-04) Restoring the density-gradient expansion for exchange in solids and surfaces. Phys. Rev. Lett. 100, pp. 136406. External Links: Document, Link Cited by: §IV.1, §IV.4.
- [25] (2020-01) Wannier90 as a community code: new features and applications. J. Phys.: Condens. Matter 32 (16), pp. 165902. External Links: Document, Link Cited by: §IV.2.
- [26] (2016) EPW: electron–phonon coupling, transport and superconducting properties using maximally localized Wannier functions. Comput. Phys. Commun. 209, pp. 116 – 133. External Links: Document Cited by: §IV.3.
@@
- [27] (2021-10) First-principles predictions of Hall and drift mobilities in semiconductors. Phys. Rev. Res. 3, pp. 043022. External Links: Document, Link Cited by: §II.2.
- [28] (2020) ABINIT: overview, and focus on selected capabilities. J. Chem. Phys. 152, pp. 124102. External Links: Document Cited by: §IV.3, §IV.3.
- [29] (2019-05) Many-body perturbation theory calculations using the yambo code. J. Phys.: Condens. Mat. 31 (32), pp. 325902. External Links: Document, Link Cited by: §IV.2.
@@
- [30] (2021) Phonon thermal transport in diamond and lonsdaleite: a comparative study of empirical potentials. Diam. Relat. Mater. 120, pp. 108618. External Links: Document, Link Cited by: §I.
@@
- [31] (2016-05) Single crystal diamond wafers for high power electronics. Diam. Relat. Mater. 65 (SI), pp. 168–175. External Links: Document, Link Cited by: §I.
@@
- [32] (2022-JUL 22) High ambipolar mobility in cubic boron arsenide. Science 377 (6604), pp. 437–440. External Links: Document, ISSN 0036-8075 Cited by: §II.1.
- [33] (2001-12) Maximally localized Wannier functions for entangled energy bands. Phys. Rev. B 65, pp. 035109. External Links: Document, Link Cited by: §IV.2.
@@
- [34] (2018) Unusual high thermal conductivity in boron arsenide bulk crystals. Science 361 (6402), pp. 582–585. External Links: Document, Link Cited by: §II.1.
- [35] (2018) The PseudoDojo: training and grading a 85 element optimized norm-conserving pseudopotential table. Comput. Phys. Commun. 226, pp. 39–54. External Links: ISSN 0010-4655, Document, Link Cited by: §IV.1.
@@
- [36] (2008) Diamond as an electronic material. Mater. Today 11 (1-2), pp. 22–28. External Links: Document, Link Cited by: §I.
@@
- [37] (2025) Synthesis of bulk hexagonal diamond. Nature (London) 644 (8076), pp. 370–375. External Links: Link, Document Cited by: §I.
@@
- [38] (2024-10) First-principles investigation of microscopic mechanisms underlying hole mobilities in diamond, silicon, and germanium. Phys. Rev. B 110, pp. 155203. External Links: Document, Link Cited by: §II.2, §II.3.
@@
- [39] (2022-JUL 22) High ambipolar mobility in cubic boron arsenide revealed by transient reflectivity microscopy. Science 377 (6604), pp. 433–436. External Links: Document, ISSN 0036-8075 Cited by: §II.1.
@@
- [40] (2025-09) Thermal properties of hexagonal diamond: machine learning potential and molecular dynamics study. Phys. Rev. Mater. 9, pp. 094603. External Links: Document, Link Cited by: §I, §II.1.
@@
- [41] (2024) Pure hexagonal diamond with symmetry-doping properties. ACS Appl. Mater. Interfaces 16 (39), pp. 52674–52681. External Links: Document, Link Cited by: §II.1.